# RadioDiff: An Effective Generative Diffusion Model for Sampling-Free Dynamic Radio Map Construction

### v2

@@ I Introduction
In wireless networks, pathloss quantifies the attenuation of signal strength between a pair of sender and receiver caused by free-space propagation loss and interactions of radio waves with obstacles [1, 2], which is critical to be measured for wireless resource management [3, 4, 5]. Traditionally, the pathloss measurement usually depends on pilot transmission and signal processing [6]. However, the dramatic increase of network nodes and antennas has led to the challenge of estimating high-dimensional channels [7], resulting in significant costs in training and feedback overhead, as well as signal processing complexity [8]. This issue further deteriorates in high-mobility scenarios with short channel coherence times for low-latency applications [9, 10]. Meanwhile, the upcoming 6G networks will introduce a large variety of node types, including passive equipment such as intelligent reflective surfaces (IRS) [11], which cannot actively transmit pilots or engage in digital signal processing to measure the pathloss.
@@
To address these challenges, numerous researchers have exploited neural networks (NNs) with rapid inference capabilities to facilitate RM construction. The most notable pioneering work is the RadioUnet that leverages the U-Net, which is a classical architecture for image-to-image tasks, for RM construction [17]. Although some NN-based methods have demonstrated better performance than traditional SPM-based methods in RM construction, their performance, especially in details construction for the dynamic RM (DRM) with multiple dynamic obstacles [^1] It is important to note that in this paper, we construct the RM based on a snapshot of the environment, without considering the Doppler effect caused by vehicle motion on the pathloss. Consequently, the main difference between DRM and SRM lies in the snapshot time interval and whether the influence of small-scale obstacles on the propagation of electromagnetic rays is considered., is poor. Therefore, the majority of NN-based methods focus on RM construction in static environments, neglecting the exploration of RM construction in dynamic environments [17, 18, 19, 20]. This is primarily attributed to two challenges: the heterogeneous propagation characteristics of electromagnetic waves and the complexity of RM texture features. The RM construction in static environments typically only concentrates on the influence of buildings on electromagnetic ray propagation, since it can be assumed that all objects affect electromagnetic rays in the same manner. In the context of DRM construction, the impact of moving objects, such as vehicles, on electromagnetic rays must be taken into account. As shown in Fig. LABEL:fig-example, the influence of vehicles on electromagnetic ray propagation markedly differs from that of buildings. Different from static obstacles with large shapes and high heights, which can completely block direct electromagnetic rays from the base station (BS) to their surfaces, dynamic obstacles such as vehicles, due to their low height and small size, cannot entirely obstruct the electromagnetic signal. This partial blockage results in a reduction of the pathloss rather than a complete obstruction, thereby increasing the complexity of RM construction and the texture features than SRM. Moreover, existing NN-based methods mainly train the NN through supervised learning in a discriminative style, whose objective is to minimize the mean square error (MSE) between the predicted RM and the ground truth. However, it has been demonstrated in [21, 22] that while the widely used MSE metric in discriminative training methods enhances the convergence speed of NN training, it degrades the NN’s ability to capture the elaborately detailed features of the data, resulting in blurred edges. Unfortunately, as shown in Fig. LABEL:fig-example RM, especially the DRM, typically has more detailed texture features, where the supervised learning training methods based on MSE struggle for NNs to learn to extract these detailed features.
@@
The fundamental reason for the above challenges is the misalignment between the generative problem attributes of the RM construction and the discriminant methodology. RM construction shows the characteristics of generative problems in both data features and NN training methods, while the existing sample-free NN-based methods use discriminant methods to construct RM, which inevitably limits the construction performance of RM. Specifically, in terms of data features, neither the value nor location of the elements of the pathloss to be predicted exist in the input environmental data, thus NN is required to generate pathloss from the raw data. Moreover, since the elements in the RM are not discrete values, it is almost impossible to construct the RM by dividing a finite number of hyperplanes to classify the elements in the environmental data as a discriminative problem [23]. Therefore, the use of discriminant training methods to generate multiple hyperplanes in the latent space using NN to predict pathloss will inevitably lead to poor performance. Furthermore, the self-supervised training method that uses data which is partially masked as input to train NN to predict masked information is mainly used in the training of generative models [24]. In the context of RM construction, the data representing environmental information, can be regarded as the RM whose pathloss is masked, and the NN-based RM construction methods use these masked RM data to recover the pathloss, which is a self-supervision training method. Based on the above analysis, RM construction is a generative problem, so from the perspective of the alignment of the problem and the methodology, a generative method should be used to effectively construct the RM. Moreover, the generative training methods enable NN training both using generative loss and MSE loss, where the NN’s capacity to extract intricate texture features of RM can be enhanced. Generative adversarial networks (GANs) have been extensively studied in RM reconstruction; however, their potential is often hindered by extreme instability during training. In contrast, the diffusion model offers a novel solution to the RM reconstruction problem. Geographic maps and RM contain abundant high-frequency sharp edges. Directly feeding these into NN-based models can result in erratic predictions, noise, and artifacts. Diffusion models, however, demonstrate superior capability in capturing this edge information and predicting conditions. Therefore, to achieve high-performance RM construction by taking advantage of the generative NN training method, in this paper, the RM construction problem is modeled as a conditional generative problem, where a generative diffusion-based method, named RadioDiff, is proposed to construct the RM effectively  [^2] In this paper, the terms “diffusion model” and “denoise diffusion model” are used interchangeably, both referring to methods for generating the required data from noise input. The diffusion process involves diffusing the data into noise, while the denoising process reconstructs the data from this noisy input.. The main contributions of this paper are as follows.
@@ II-B Diffusion Model
The diffusion model is a category of generative models based on Markov chains, that progressively restore data through a learned denoising process. This model has emerged as a strong competitor to the GAN in various generative tasks, such as computer vision [25], and natural language processing [21]. Moreover, diffusion models exhibit significant potential in perception tasks, including image segmentation, object detection, and model-based reinforcement learning (RL) [26]. In the diffusion model, there are two procedures which are the forward diffusion procedure where the raw data is diffused into noise, and the reversed denoise procedure where an NN is used to remove the noise adding to the raw data, thus generating raw data from noise.

### v3

@@ I Introduction
In wireless networks, pathloss quantifies the attenuation of signal strength between a pair of sender and receiver caused by free-space propagation loss and interactions of radio waves with obstacles [1, 2, 3], which is critical to be measured for wireless resource management [4, 5, 6]. Traditionally, the pathloss measurement usually depends on pilot transmission and signal processing [7]. However, the dramatic increase of network nodes and antennas has led to the challenge of estimating high-dimensional channels [8], resulting in significant costs in training and feedback overhead, as well as signal processing complexity [9]. This issue further deteriorates in high-mobility scenarios with short channel coherence times for low-latency applications [10, 11]. Meanwhile, the upcoming 6G networks will introduce a large variety of node types, including passive equipment such as intelligent reflective surfaces (IRS) [12], which cannot actively transmit pilots or engage in digital signal processing to measure the pathloss.
@@
[…]
@@
The fundamental reason for the above challenges is the misalignment between the generative problem attributes of the RM construction and the discriminant methodology. RM construction shows the characteristics of generative problems in both data features and NN training methods, while the existing sample-free NN-based methods use discriminant methods to construct RM, which inevitably limits the construction performance of RM. Specifically, in terms of data features, neither the value nor location of the elements of the pathloss to be predicted exist in the input environmental data, thus NN is required to generate pathloss from the raw data. Moreover, since the elements in the RM are not discrete values, it is almost impossible to construct the RM by dividing a finite number of hyperplanes to classify the elements in the environmental data as a discriminative problem [24, 25]. Therefore, the use of discriminant training methods to generate multiple hyperplanes in the latent space using NN to predict pathloss will inevitably lead to poor performance. Furthermore, the self-supervised training method that uses data which is partially masked as input to train NN to predict masked information is mainly used in the training of generative models [26]. In the context of RM construction, the data representing environmental information, can be regarded as the RM whose pathloss is masked, and the NN-based RM construction methods use these masked RM data to recover the pathloss, which is a self-supervision training method. Based on the above analysis, RM construction is a generative problem, so from the perspective of the alignment of the problem and the methodology, a generative method should be used to effectively construct the RM. Moreover, the generative training methods enable NN training both using generative loss and MSE loss, where the NN’s capacity to extract intricate texture features of RM can be enhanced. Generative adversarial networks (GANs) have been extensively studied in RM reconstruction; however, their potential is often hindered by extreme instability during training. In contrast, the diffusion model offers a novel solution to the RM reconstruction problem. Geographic maps and RM contain abundant high-frequency sharp edges. Directly feeding these into NN-based models can result in erratic predictions, noise, and artifacts. Diffusion models, however, demonstrate superior capability in capturing this edge information and predicting conditions. Therefore, to achieve high-performance RM construction by taking advantage of the generative NN training method, in this paper, the RM construction problem is modeled as a conditional generative problem, where a generative diffusion-based method, named RadioDiff, is proposed to construct the RM effectively  [^2] In this paper, the terms “diffusion model” and “denoise diffusion model” are used interchangeably, both referring to methods for generating the required data from noise input. The diffusion process involves diffusing the data into noise, while the denoising process reconstructs the data from this noisy input.. The main contributions of this paper are as follows.
@@ II-B Diffusion Model
The diffusion model is a category of generative models based on Markov chains, that progressively restore data through a learned denoising process. This model has emerged as a strong competitor to the GAN in various generative tasks, such as computer vision [27], and natural language processing [22]. Moreover, diffusion models exhibit significant potential in perception tasks, including image segmentation, object detection, and model-based reinforcement learning (RL) [28]. For example, in [29], the authors effectively utilized diffusion model-based soft actor-critic algorithms for optimal contract design. Besides, the authors in [30] innovatively leveraged diffusion model-based deep deterministic policy gradient algorithms for optimal Stackelberg game solutions. In the diffusion model, there are two procedures which are the forward diffusion procedure where the raw data is diffused into noise, and the reversed denoise procedure where an NN is used to remove the noise adding to the raw data, thus generating raw data from noise.